OpenStudy (anonymous):

elimination method 0.3x-0.2y=4 0.5x+0.5y=-45/17

OpenStudy (anonymous):

ick who gave this question with a 17 in the denominator?

OpenStudy (anonymous):

must be tired. start with \[3x-2y=40\] \[x+y=-\frac{90}{17}\]

myininaya (myininaya):

(1) \[\frac{3}{10}x-\frac{2}{10}y=4\] (2) \[\frac{5}{10}x+\frac{5}{10}y=\frac{-45}{17}\] multiply (1) by 10 multiply (2) by 10(17) \[3x-2y=40\] \[5(17)x+5(17)y=-45(10)\]

myininaya (myininaya):

(1) \[3x-2y=40\] (2) \[5(17)x+5(17)y=-45(10) \] divide equation (2) by 5

myininaya (myininaya):

(1) \[3x-2y=40\] (2) \[17x+17y=-45(2)\] so we have equation (2) is \[17x+17y=-90\]

OpenStudy (anonymous):

multiply second equation by 2 get \[3x-2y=40\] \[2x+2y=-\frac{180}{17}\] add and get \[5x=40-\frac{180}{17}=\frac{500}{17}\] so \[x=\frac{100}{17}\]

myininaya (myininaya):

now are two equations look pretty right?

myininaya (myininaya):

ok but satellite was much quicker so use his process

OpenStudy (anonymous):

(blush)

OpenStudy (anonymous):

or you can use the decimal w/c is the orig eq 0.3x-0.2y=4 mult by 0.5 0.15x-0.10y= 2 0.5x+0.5y=-45/17 mult by 0.2 +(0.10x+0.10y=-9/17) add -------------------- 0.25x +0 = 25/17 x=25/(17)(.25) x=100/17 ans now sub this to 0.5x+0.5y=-45/17 0.5(100/17)+0.5y=-45/17 y=-11.176 ans

OpenStudy (anonymous):

good luck stacy